# A Survey of Model Architectures in Information Retrieval

### v2

@@ 4 Learning-to-Rank Model Architectures
Despite their success, traditional LTR models have a fundamental ceiling. Their reliance on handcrafted features means they are not end-to-end and, more critically, they struggle to bridge the lexical gap — the difference between the words in a query and the semantically related words in a relevant document. Their understanding is based on pre-defined statistical signals, not the underlying meaning of the text. This limitation created a clear need for a new class of models capable of learning semantic representations directly from raw text, setting the stage for the rise of neural ranking (§˜5).
@@ 5 Neural Ranking Models
Neural ranking models emerged to directly address the key limitations of LTR (§˜4). By learning semantic representations directly from raw text, they could automatically bridge the lexical gap — for instance, recognizing that a query for “computer” is conceptually related to a document about a “PC” without relying on term overlap. This end-to-end approach simultaneously eliminated the laborious process of manual feature engineering, shifting the paradigm from engineering statistical signals to learning semantic patterns from data. [^1] For the sake of paper structure, in this section we focus on neural information retrieval which cover retrieval models based on neural networks prior to pre-trained transformers. We kindly refer more details to the dedicated surveys (Onal et al., 2018; Mitra et al., 2018; Xu et al., 2018).
@@
Depending on how queries interact with documents during network processing, neural ranking models can be roughly divided into representation-based models and interaction-based models (Guo et al., 2016a). This division reflects a fundamental trade-off between efficiency and matching depth. Representation-based models pre-encode documents into vectors offline, enabling highly efficient retrieval suitable for first-pass ranking. In contrast, interaction-based models process the query and document together, allowing for deeper and more precise matching at a higher computational cost, making them ideal for reranking a smaller set of candidates.
Representation-based models.
@@
This genre of models can be regarded as an extension of vector space models (§˜3), which independently encode queries and documents into a shared latent vector space, where relevance is determiend through simple comparison functions such as cosine similarity or dot product, as illustrated in Fig.˜2(a). This approach maintains clear separation between query and document processing, with no interaction occurring during the encoding procedure. The core architectural challenge is designing an encoder network that transforms a variable-length sequence of term embeddings into a single, fixed-size semantic vector.
@@
The Deep Structured Semantic Model (DSSM, Huang et al., 2013; Gao et al., 2014) is an early example. It utilizes word hashing (a technique to manage large vocabularies by grouping words into a smaller number of hash buckets) and multilayer perceptrons (MLPs) to independently encode term vectors of queries and documents, enabling the computation of ranking scores based on the cosine similarity of their embeddings. Later works modify DSSM’s encoder network to better capture richer semantic and contextual information. Convolutional DSSM(C-DSSM, Shen et al., 2014a) leverages a CNN architecture. Specifically, it applies 1D convolutions over the sequence of word embeddings, allowing the model to learn representations for n-grams and local phrases. A max-pooling layer then selects the most salient local features to form the final document vector. Another variant of DSSM replaces MLPs with recurrent layers such as Long Short-Term Memory (LSTM) network (Hochreiter & Schmidhuber, 1997; Palangi et al., 2016; Wan et al., 2016; Cohen & Croft, 2016) or tree-structured networks (Tai et al., 2015). The LSTM processes the text sequentially, and its recurrent nature allows it to capture word order and long-range dependencies across the entire text, with the final hidden state often used as the comprehensive representation for the query or document.
@@
These models excel in scenarios requiring global semantic understanding and offer significant computational advantages through their ability to pre-compute document representations offline (Guo et al., 2019). However, representation-focused approaches face inherent limitations due to their reliance on fixed-size embedding vectors, which can struggle to capture all relevant information from the original text and may not effectively handle precise lexical matching requirements. These limitations are the focus of interaction-based models.
Interaction-based models.
@@
Different from representation-based models, interaction-based models (Fig.˜2(b)) process queries and documents jointly through neural networks. Instead of compressing each text into a single vector, they first build a detailed, low-level interaction representation between the query and the document, and then use neural networks to learn hierarchical matching patterns from this representation. The model’s output is typically a scalar relevance score of the input query-document pair. Various network architectures have been proposed under this paradigm. MatchPyramid (Pang et al., 2016) employs CNN over the interaction matrix between query and document terms. The interaction matrix is treated as an image, allowing the CNN with its 2D filters to capture local matching patterns (e.g., phrases or bigrams matching) through convolution and pooling operations (Hu et al., 2014).
@@
Building upon the concept of interaction-focused models, Guo et al. (2016a) highlight the importance of exact term matches in neural ranking models and proposed the Deep Relevance Matching Model (DRMM). Rather than a single interaction matrix, DRMM creates a matching histogram for each query term. This histogram discretizes the similarity scores against all document terms into bins, effectively capturing the distribution of matching signals (e.g., how many terms in the document are an exact match, a strong semantic match, or a weak match to a given query term). An MLP then learns the relevance contribution from these histogram features.
@@
In line with these works, the interaction matrix-based approach have been explored for short text matching (Lu & Li, 2013; Yin et al., 2016; Yang et al., 2016) as well as long document ranking (Mitra et al., 2017; Hui et al., 2017; 2018, inter alia). Multi-Perspective CNN approaches compare sentences via diverse pooling functions and filter widths to capture multiple perspectives between texts He et al. (2015). Attention-based methods like aNMM (Yang et al., 2016) perform soft-attention alignment by measuring pairwise word similarity between each word in query and answer, while attentive pooling models utilize this alignment to learn weighted representations that depend on each other. Term adjacency and positional information represent another important dimension of interaction modeling. Models such as MatchPyramid, PACRR, and ConvKNRM capture term adjacency patterns and position-dependent interactions (Pang et al., 2016; Hui et al., 2017; Dai et al., 2018).
The interaction functions in these models can be categorized as either non-parametric (using traditional similarity measures like cosine similarity, dot product, or binary indicators) or parametric (learning similarity functions from data through neural networks) (Dong et al., 2022). While interaction-focused models require one forward pass through the entire model for each potentially relevant document, making them computationally more expensive than representation-focused approaches, they typically achieve superior ranking quality due to their ability to capture fine-grained matching signals. We list some representative works in Table˜2 and direct readers to these works for architectural details.
Hybrid Models.
@@
This hybrid architecture demonstrates that combining distributed representations with traditional local representations is favorable, with the combined approach significantly outperforming either neural network individually. More recent hybrid approaches have focused on reducing computational costs while maintaining effectiveness, with some models incorporating cached token-level representations to enable faster query-document interactions when document representations are pre-computed (Wrzalik & Krechel, 2020). The success of hybrid models has established that interaction-based and representation-based approaches can be effectively combined for further improvements in ranking performance (Liu et al., 2018).
@@
DESM Nalisnick et al. (2016a) Representation-based MLP Randomly Initialized
@@
aNMM (Yang et al., 2016) Interaction-based Attention Word2Vec
DESM (Nalisnick et al., 2016b) Interaction-based MLP Word2Vec
@@
In addition to the development of network architecture, pre-trained embeddings (Salakhutdinov & Hinton, 2009; Mikolov, 2013; Pennington et al., 2014; Le & Mikolov, 2014) provide semantic-based term representations to enable neural ranking models to focus on learning relevance matching patterns, improving training convergence and retrieval performance on both representation-based and interaction-based models (Levy et al., 2015). Both GloVe (Pennington et al., 2014) and Word2Vec (Mikolov, 2013) learn dense vector representations for each vocabulary term from large-scale text corpora. By initializing the embedding layer with these pre-trained vectors, models start with a strong semantic foundation, which proved crucial for performance, especially on smaller training datasets (Guo et al., 2016b). Interaction-based models with cross-lingual word embeddings (Joulin et al., 2018) for cross-lingual reranking have also been explored (Yu & Allan, 2020). Table˜2 shows a list of neural ranking models and backbone architectures. Researchers have explored different backbone neural network architectures in this era, including Convolutional Neural Network (CNN, LeCun et al., 1989), Long Short Term Memory (LSTM, Hochreiter & Schmidhuber, 1997) and kernel methods (Vert et al., 2004; Chang et al., 2010; Xiong et al., 2017).
@@
Notably, a line of research explores integrating kernel methods with the Transformer architecture (Vaswani et al., 2017). The main distinction between this line of research and the models discussed in §˜6 is that the transformer modules here are not pre-trained on large-scale corpora like Wikipedia and C4 (Devlin et al., 2019; Raffel et al., 2020). We consider this line of research as an intersection between neural ranking models (§˜5) and retrieval with pre-trained transformers (§˜6). TK (Hofstätter et al., 2020c) uses a shallow transformer neural network (up to 3 layers) to encode the query $\mathcal{Q}$ and document $\mathcal{D}$ separately. After encoding, the contextualized representations are input to an interaction module inspired by K-NRM, where RBF kernels are used to create soft-match features from the contextualized embeddings. This fusion of a transformer encoder with a kernel-based interaction mechanism allowed the model to achieve better performance-efficiency trade-off compared to BERT-based reranker (Nogueira et al., 2019b). The main bottleneck of applying transformer architectures to long document reranking is $O(n^{2})$ time complexity, where $n$ denotes the document length. TKL (Hofstätter et al., 2020b) further improves upon TK with a local attention mechanism and leads to performance improvement on long document ranking.
@@
The neural ranking models described above, particularly later developments like TK and TKL, demonstrated the potential of the transformer’s attention mechanism for modeling relevance. However, the true paradigm shift occurred when the IR community moved from using these architectures trained from scratch to leveraging massive, pre-trained transformer models like BERT (Devlin et al., 2019) and its variants (Liu, 2019; Sun et al., 2019; Lan et al., 2020; Beltagy et al., 2020). This marked a fundamental change in approach: instead of designing novel, task-specific network backbones (e.g., CNNs, LSTMs) on top of static word embeddings, research shifted to fine-tuning a single, powerful, and deeply contextualized architecture for IR tasks. This new foundation did not eliminate the core architectural trade-offs but rather recast them in a more powerful form, leading to the development of cross-encoder rerankers and bi-encoder retrievers, which we explore next.
@@ 6 IR with Pre-trained Transformers
(b) Learned dense retrieval
@@
Figure 3: Illustration on transformer-based retrieval and reranking models. Yellow boxes indicate pretrained Transformers (e.g., BERT).
@@
BERT (Devlin et al., 2019) revolutionized research in both natural language processing (NLP) and information retrieval (IR). Its success is largely attributed to two key factors: (1) the Multi-Head Attention (MHA) architecture (Vaswani et al., 2017), which enables high-dimensional, contextualized token representations; and (2) large-scale pre-training, which equips BERT with the ability to capture rich semantics and world knowledge. The expressive power of BERT has been extensively analyzed in prior work, e.g., (Rogers et al., 2020; Tenney et al., 2019; Clark, 2019).
@@
Before discussing specific IR modeling architectures, it is useful to understand the high-level architectural families of transformers, as they have distinct implications for IR tasks:
@@
  • Encoder-only models (e.g., BERT (Devlin et al., 2019), RoBERTa (Liu, 2019)) use a bidirectional self-attention mechanism, allowing each token’s representation to be informed by the entire input sequence (both left and right context). This deep contextual understanding makes them naturally suited for representation-focused tasks. In IR, they have been the workhorses for building powerful bi-encoder retrievers and cross-encoder rerankers.
@@
  • Decoder-only models (e.g., proprietary GPT series (Radford et al., 2019; Brown et al., 2020; OpenAI et al., 2024), open-weight Llama (Touvron et al., 2023), Mistral (Jiang et al., 2023)) use a unidirectional (causal) self-attention mechanism, where each token can only attend to previous tokens in the sequence. This architecture is optimized for next-token prediction and, by extension, text generation. Their application to representation tasks like retrieval is less direct and often requires architectural adaptations to create meaningful summary vectors from their unidirectional hidden states.
@@
  • Encoder-decoder models (e.g., T5 (Raffel et al., 2020), BART (Lewis et al., 2020a)) combine both architectures. The encoder processes the input sequence bidirectionally to create a rich representation, which then conditions the decoder to generate an output sequence autoregressively. This "sequence-to-sequence" design makes them highly versatile. In IR, they can be framed as rerankers (generating a "relevant" or "irrelevant" token), or as generative retrievers that directly generate document identifiers.
@@
This section discusses IR architectures based on pre-trained transformers, with a focus on BERT-type encoder models, which is to be distinct from encoder-decoder models and decoder-only models covered in §˜7. We structure our review around the fundamental architectural trade-off between interaction depth and computational efficiency. We begin with models that perform deep, full interaction (cross-encoders), which are powerful but slow, making them ideal for reranking. We then move to separable architectures (bi-encoders) that prioritize efficiency for first-stage retrieval, exploring their dense, sparse, and multi-vector variants. Finally, we discuss orthogonal improvements such as hybrid models, continual training, and interpretability.
Deep Interaction Models: The Cross-Encoder for Reranking.

### v3

@@ 4.3 Orthogonal Directions
Despite their success, traditional LTR models have a fundamental ceiling. Their reliance on handcrafted features prevents them from operating in an end-to-end manner and, more importantly, limits their ability to bridge the lexical gap — the difference between the words in a query and the semantically related words in a relevant document. Their understanding is based on pre-defined statistical signals, not the underlying meaning of the text. This limitation created a clear need for a new class of models capable of learning semantic representations directly from raw text, setting the stage for the rise of neural ranking (Section˜5).
@@ 5 Neural Ranking Models
Neural ranking models emerged to directly address the key limitations of LTR (Section˜4). By learning semantic representations directly from raw text, they could automatically bridge the lexical gap — for instance, recognizing that a query for “computer” is conceptually related to a document about a “PC” without relying on term overlap. This end-to-end approach simultaneously eliminated the laborious process of manual feature engineering, shifting the paradigm from engineering statistical signals to learning semantic patterns from data. [^1] For structural clarity, this section focuses on neural information retrieval, covering neural network–based retrieval models developed prior to the advent of pre-trained transformers. For a more comprehensive treatment, we refer readers to the dedicated surveys (Onal et al., 2018; Mitra et al., 2018; Xu et al., 2018).
@@
Depending on how queries interact with documents during network processing, neural ranking models can be roughly divided into representation-based models and interaction-based models (Guo et al., 2016a). This division reflects a fundamental tradeoff between efficiency and matching depth. Representation-based models pre-encode documents into vectors offline, enabling highly efficient retrieval suitable for first-pass ranking. In contrast, interaction-based models process the query and document together, allowing for deeper and more precise matching at a higher computational cost, making them ideal for reranking a smaller set of candidates.
5.1 Representation-Based Models
@@
This genre of models can be regarded as an extension of Vector Space Models (Section˜3), which independently encode queries and documents into a shared latent vector space, where relevance is determined through simple comparison functions such as cosine similarity or dot product, as illustrated in Figure˜2(a). This approach maintains clear separation between query and document processing, with no interaction occurring during the encoding procedure. The core architectural challenge is designing an encoder network that transforms a variable-length sequence of term embeddings into a single, fixed-size semantic vector.
@@
The Deep Structured Semantic Model (DSSM, Huang et al., 2013; Gao et al., 2014) is an early example. It utilizes word hashing (a technique to manage large vocabularies by grouping words into a smaller number of hash buckets) and multilayer perceptrons (MLPs) to independently encode term vectors of queries and documents, enabling the computation of ranking scores based on the cosine similarity of their embeddings. Later works modify DSSM’s encoder network to better capture richer semantic and contextual information. Convolutional DSSM(C-DSSM, Shen et al., 2014a) leverages a convolutional neural network (CNN) architecture. Specifically, it applies 1D convolutions over the sequence of word embeddings, allowing the model to learn representations for n-grams and local phrases. A max-pooling layer then selects the most salient local features to form the final document vector. Another variant of DSSM replaces MLPs with recurrent layers such as Long Short-Term Memory (LSTM) network (Hochreiter & Schmidhuber, 1997; Palangi et al., 2016; Wan et al., 2016; Cohen & Croft, 2016) or tree-structured networks (Tai et al., 2015). The LSTM processes the text sequentially, and its recurrent nature allows it to capture word order and long-range dependencies across the entire text, with the final hidden state often used as the comprehensive representation for the query or document. Based on the assumption of documents’ hierarchical structure, Yang et al. (2016b); Song et al. (2018); Zhu et al. (2019) use attention (Bahdanau et al., 2014) to model token, phrase and sentence representations for enhanced document/passage representations.
In line with these works, the NLP community has also extensively investigated passage/document representations. Le & Mikolov (2014) proposed Paragraph Vectors (Doc2Vec), an unsupervised algorithm that learns fixed-length feature representations from variable-length pieces of texts, which is based on single-hidden-layer neural network. Kim (2014) studied CNN for sentence representations, while Wieting et al. (2016) proposed to use LSTM network. Arora et al. (2017) reported a weighted average of pre-trained word embeddings fine-tuned with unsupervised random walk algorithm can outperform more complex neural networks on the sentence similarity task. However, weighted averaging word embeddings ignores the word order, which fails to capture the rich, contextual information in longer, more complex documents.
@@
To summarize, representation-based models excel in scenarios requiring global semantic understanding and offer significant computational advantages through their ability to pre-compute document representations offline (Guo et al., 2020). However, these approaches also face inherent limitations due to their reliance on fixed-size embedding vectors, which can struggle to capture all relevant information from the original text and may not effectively handle precise lexical matching requirements. These limitations are the focus of interaction-based models.
5.2 Interaction-Based Models
@@
Different from representation-based models, interaction-based models (Figure˜2(b)) process queries and documents jointly through neural networks. Instead of compressing each text into a single vector, they first build a detailed, low-level interaction representation between the query and the document, and then use neural networks to learn hierarchical matching patterns from this representation. The model’s output is typically a scalar relevance score of the input query-document pair. Various network architectures have been proposed under this paradigm. MatchPyramid (Pang et al., 2016) employs CNN over the interaction matrix between query and document terms. The interaction matrix is treated as an image, allowing the CNN with its 2D filters to capture local matching patterns (e.g., phrases or bigrams matching) through convolution and pooling operations (Hu et al., 2014).
@@
Building upon the concept of interaction-focused models, Guo et al. (2016a) highlighted the importance of exact term matches in neural ranking models and proposed the Deep Relevance Matching Model (DRMM). Rather than a single interaction matrix, DRMM creates a matching histogram for each query term. This histogram discretizes the similarity scores against all document terms into bins, effectively capturing the distribution of matching signals (e.g., how many terms in the document are an exact match, a strong semantic match, or a weak match to a given query term). An MLP then learns the relevance contribution from these histogram features.
@@
In line with these works, the interaction matrix-based approach have been explored for short text matching (Lu & Li, 2013; Yin et al., 2016; Yang et al., 2016a) as well as long document ranking (Mitra et al., 2017; Hui et al., 2017; 2018, inter alia). Multi-Perspective CNN approaches compare sentences via diverse pooling functions and filter widths to capture multiple perspectives between texts He et al. (2015). aNMM (Yang et al., 2016a), as an example of attention-based methods, computes passage terms’ attention weights over query terms using a query attention network and achieves performance improvement compared to CNN-based baseline (Severyn & Moschitti, 2015). Term adjacency and positional information represent another important dimension of interaction modeling. Models such as MatchPyramid, PACRR, and ConvKNRM capture term adjacency patterns and position-dependent interactions (Pang et al., 2016; Hui et al., 2017; Dai et al., 2018).
The interaction functions in these models can be categorized as either non-parametric (using traditional similarity measures like cosine similarity, dot product, or binary indicators) or parametric (learning similarity functions from data through neural networks) (Dong et al., 2022b). While interaction-focused models require one forward pass through the entire model for each potentially relevant document, making them computationally more expensive than representation-focused approaches, they typically achieve superior ranking quality due to their ability to capture fine-grained matching signals. We list some representative works in Table˜2 and direct readers to these works for architectural details.
5.3 Hybrid Models
@@
Different from the metric learning theme of representation-based models, a line of works formulates the ranking problem as a classification problem (commonly referred to as Extreme Label Classification, or XMC), where the input is the query, and the output is a probability distribution over the corpus, where each document is a unique “class” or “label”. Instead of optimizing the similarity between query and document representations, XMC models aim to predict the correct subset of relevant document IDs (Prabhu & Varma, 2014; Jain et al., 2016; Liu et al., 2017; Jain et al., 2019). In the inference time, XMC methods use tree-based hierarchies or cluster-based sampling to quickly narrow down the search path to the likely labels without scanning every candidates (Prabhu et al., 2018; You et al., 2019). AttentionXML (You et al., 2019) uses an attention mechanism to focus on specific parts of the input text that are most relevant to the label’s semantic meaning, and thus can be considered a hybrid model. We refer readers to (Dasgupta et al., 2023) for a comprehensive review of XMC methods.
This hybrid architecture demonstrates that combining distributed representations with traditional local representations is favorable, with the combined approach significantly outperforming either neural network individually. More recent hybrid approaches have focused on reducing computational costs while maintaining effectiveness, with some models incorporating cached token-level representations to enable faster query-document interactions when document representations are pre-computed (Wrzalik & Krechel, 2021). The success of hybrid models has established that interaction-based and representation-based approaches can be effectively combined for further improvements in ranking performance (Liu et al., 2018).
@@
DESM Nalisnick et al. (2016) Representation-based MLP Randomly Initialized
@@
aNMM (Yang et al., 2016a) Interaction-based Attention Word2Vec
DESM (Nalisnick et al., 2016) Interaction-based MLP Word2Vec
@@ 5.4 Orthogonal Directions
In addition to the development of network architecture, pre-trained embeddings (Salakhutdinov & Hinton, 2009; Mikolov et al., 2013; Pennington et al., 2014; Le & Mikolov, 2014) provide semantic-based term representations to enable neural ranking models to focus on learning relevance matching patterns, improving training convergence and retrieval performance on both representation-based and interaction-based models (Levy et al., 2015). Both GloVe (Pennington et al., 2014) and Word2Vec (Mikolov et al., 2013) learn dense vector representations for each vocabulary term from large-scale text corpora. By initializing the embedding layer with these pre-trained vectors, models start with a strong semantic foundation, which proved crucial for performance, especially on smaller training datasets (Guo et al., 2016b). Interaction-based models with cross-lingual word embeddings (Joulin et al., 2018) for cross-lingual reranking have also been explored (Yu & Allan, 2020). Table˜2 shows a list of neural ranking models and backbone architectures. Researchers have explored different backbone neural network architectures in this era, including Convolutional Neural Network (CNN, LeCun et al., 1989), Long Short Term Memory (LSTM, Hochreiter & Schmidhuber, 1997) and kernel methods (Vert et al., 2004; Chang et al., 2010; Xiong et al., 2017).
@@
Notably, a line of research explores integrating kernel methods with the Transformer architecture (Vaswani et al., 2017). The main distinction between this line of research and the models discussed in Section˜6 is that the transformer modules here are not pre-trained on large-scale corpora like Wikipedia and C4 (Devlin et al., 2019; Raffel et al., 2020). We consider this line of research as an intersection between neural ranking models (Section˜5) and retrieval with pre-trained transformers (Section˜6). TK (Hofstätter et al., 2020c) uses a shallow transformer neural network (up to 3 layers) to encode the query $\mathcal{Q}$ and document $\mathcal{D}$ separately. After encoding, the contextualized representations are input to an interaction module inspired by K-NRM, where RBF kernels are used to create soft-match features from the contextualized embeddings. This fusion of a transformer encoder with a kernel-based interaction mechanism allowed the model to achieve better performance-efficiency tradeoff compared to BERT-based reranker (Nogueira et al., 2019b). The main bottleneck of applying transformer architectures to long document reranking is $O(n^{2})$ time complexity, where $n$ denotes the document length. TKL (Hofstätter et al., 2020b) further improves upon TK with a local attention mechanism and leads to performance improvement on long document ranking.
@@
The neural ranking models described above, particularly later developments like TK and TKL, demonstrated the potential of the transformer’s attention mechanism for modeling relevance. However, the true paradigm shift occurred when the IR community moved from using these architectures trained from scratch to leveraging massive, pre-trained transformer models like BERT (Devlin et al., 2019) and its variants (Liu et al., 2019; Sun et al., 2019; Lan et al., 2020; Beltagy et al., 2020). This marked a fundamental change in approach: instead of designing novel, task-specific network backbones (e.g., CNNs, LSTMs) on top of static word embeddings, research shifted to fine-tuning a single, powerful, and deeply contextualized architecture for IR tasks. This new foundation did not eliminate the core architectural tradeoffs but rather recast them in a more powerful form, leading to the development of cross-encoder rerankers and bi-encoder retrievers, which we explore next.
@@ 6 IR with Pre-Trained Transformers
(b) Learned dense retrieval. ($w$ are token-level scalar weights)
@@
Figure 3: Illustration on transformer-based retrieval and reranking models. Yellow boxes indicate pretrained Transformers (e.g., BERT). Query text, embeddings, and associated weights are color-coded in orange, whereas document representations are color-coded in blue.
@@
BERT (Devlin et al., 2019) revolutionized research in both natural language processing (NLP) and information retrieval (IR). Its success is largely attributed to two key factors: (1) the Multi-Head Attention (MHA) architecture (Vaswani et al., 2017), which enables high-dimensional, contextualized token representations; and (2) large-scale pre-training, which equips BERT with the ability to capture rich semantics and world knowledge. The expressive power of BERT has been extensively analyzed in prior work, e.g., (Rogers et al., 2020; Tenney et al., 2019; Clark et al., 2019).
High-level architectural families.
@@
Before discussing specific IR modeling architectures, it is useful to understand the architectural families of transformers, as they have distinct implications for IR tasks:
@@
  • Encoder-only models (e.g., BERT (Devlin et al., 2019), RoBERTa (Liu et al., 2019)) use a bidirectional self-attention mechanism, allowing each token’s representation to be informed by the entire input sequence (both left and right context). This deep contextual understanding makes them naturally suited for representation-focused tasks. In IR, they have been the workhorses for building powerful bi-encoder retrievers and cross-encoder rerankers.
@@
  • Decoder-only models (e.g., proprietary GPT series (Radford et al., 2019; Brown et al., 2020; OpenAI et al., 2024), open-weight Llama (Touvron et al., 2023), Mistral (Jiang et al., 2023a)) use a unidirectional (causal) self-attention mechanism, where each token can only attend to previous tokens in the sequence. This architecture is optimized for next-token prediction and, by extension, text generation. Their application to representation tasks like retrieval is less direct and often requires architectural adaptations to create meaningful summary vectors from their unidirectional hidden states.
@@
  • Encoder-decoder models (e.g., T5 (Raffel et al., 2020), BART (Lewis et al., 2020a)) combine both architectures. The encoder processes the input sequence bidirectionally to create a rich representation, which then conditions the decoder to generate an output sequence autoregressively. This “sequence-to-sequence” design makes them highly versatile. In IR, they can be framed as rerankers (generating a “relevant” or “irrelevant” token), or as generative retrievers that directly generate document identifiers.
@@
This section discusses IR architectures based on pre-trained transformers, with a focus on BERT-type encoder models, which is to be distinct from encoder-decoder models and decoder-only models covered in Section˜7. We structure our review around the fundamental architectural tradeoff between interaction depth and computational efficiency. These constraints necessitate two primary paradigms:
  1. 1.
    Cross-Encoder (Deep Interaction): A single model processes the concatenated query and document as one sequence, allowing every query token to interact deeply with every document token. This provides state-of-the-art ranking quality but is computationally expensive, making it suitable only for reranking.
  2. 2.
    Bi-Encoder (Separable Pre-computation): Separate encoders process the query and document independently to create fixed-size vectors. Since document vectors can be pre-computed offline, this architecture enables extremely fast similarity search suitable for first-stage retrieval.
We first discuss the crucial training strategies used to optimize these two architectures. We then detail the cross-encoder models that perform deep, full interaction, followed by the separable bi-encoder architectures that prioritize efficiency, exploring their dense, sparse, and multi-vector variants. Finally, we discuss advanced hybrid models and orthogonal improvements such as continual training and interpretability.
6.1 Training Strategies for Transformer-Based IR
Although we aim to disentangle model architectures from training strategies, the co-evolvement of these two areas is a defining pattern of this era. The architectural dichotomy (cross-encoder vs. bi-encoder) has an impactful influence the training methodology, extending the loss function categories discussed in Section˜4 into the deep learning paradigm.
Contrastive and Listwise Objectives.
The application of contrastive learning builds on the principle of the InfoNCE loss (Oord et al., 2018), which is derived from Noise-Contrastive Estimation (Gutmann & Hyvärinen, 2012). The general goal is to learn a model that distinguishes a “positive” sample from a set of “negative” samples.
The InfoNCE framework is primarily used to optimize bi-encoders in the dense retrieval setting. In this context, the relevance score $f$ is a simple similarity function (e.g., dot product) between the query and document vectors, $f(\mathcal{Q},\mathcal{D})=\mathrm{sim}(\mathbf{v}_{\mathcal{Q}},\mathbf{v}_{\mathcal{D}})$. For a query $\mathcal{Q}_{i}$, a positive document $\mathcal{D}_{i}^{+}$, and a set of negative documents $\mathcal{D}_{i}^{-}$, the bi-encoder is trained to minimize the negative log probability of correctly classifying the positive document:
$\mathcal{L}_{\mathrm{InfoNCE}}=-\frac{1}{|\mathcal{S}|}\sum\limits_{(\mathcal{Q}_{i},\mathcal{D}_{i}^{+})\in\mathcal{S}}\log\frac{\exp f_{\theta}(\mathcal{Q}_{i},\mathcal{D}_{i}^{+})}{\exp f_{\theta}(\mathcal{Q}_{i},\mathcal{D}_{i}^{+})+\sum\limits_{\mathcal{D}_{j}^{-}\in\mathcal{D}_{i}^{-}}\exp f_{\theta}(\mathcal{Q}_{i},\mathcal{D}_{j}^{-})}$
where $\mathcal{S}$ is the training set and $\mathcal{D}_{i}^{-}$ is the set of sampled negative documents.
For cross-encoders, which compute relevance over a concatenated list, the objective is also listwise but often simplified to a standard Negative Log-Likelihood (NLL) loss over the final softmax probabilities of the candidates, minimizing the distance between the predicted distribution and the ground truth relevance distribution for the entire list.
Hard Negative Mining.
A critical challenge in training bi-encoders is generating sufficiently difficult negative examples, as random sampling typically yields easy negatives that do not challenge the model effectively. This is where Hard Negative Mining (HNM) becomes essential. HNM strategies ensure that the model is exposed to challenging cases where positive and negative document features are difficult to distinguish. Key strategies include:
  1. 1.
    In-Batch Negatives (IBN): Leveraging other queries’ positive documents within the same mini-batch as negative examples for the current query. IBN provides a balance of efficiency and difficulty.
  2. 2.
    Lexical Negatives: Using documents highly ranked by a traditional sparse model (e.g., BM25) but not labeled as relevant.
  3. 3.
    Iterative Hard Negative Mining: Employing an existing dense retriever to periodically mine difficult negatives from the collection (i.e., documents that the current model mistakenly ranks highly). Seminal methods like ANCE (Xiong et al., 2020) and ADORE (Zhan et al., 2021) use this iterative approach to continually feed the model better training data.
Knowledge Distillation.
To further narrow the effectiveness gap between efficient bi-encoders and powerful cross-encoders, the community widely adopted Knowledge Distillation (KD) (Buciluă et al., 2006; Hinton et al., 2015). KD is a technique for training a smaller, efficient “student” model by transferring knowledge from a larger, more capable “teacher” model (Hinton et al., 2015; Gou et al., 2021). In IR, KD is used to create fast rerankers or retrievers that approximate the performance of slower, larger models, which is critical for production systems (Hofstätter et al., 2020a; 2021a; Chen et al., 2021; Kim et al., 2023; Xu et al., 2025c; Zhang et al., 2025b; Vera et al., 2025).
Let $f_{t}$ denote the teacher model (typically a cross-encoder) and $f_{s}$ denote the student model (typically a bi-encoder or a smaller cross-encoder). For a given query $\mathcal{Q}$ and a list of candidate texts $\mathcal{D}=\{\mathcal{D}_{1},\mathcal{D}_{2},\ldots,\mathcal{D}_{k}\}$, we first compute relevance scores (logits) from both models:
$\mathbf{z}_{t}$ $=[f_{t}(\mathcal{Q},\mathcal{D}_{1}),f_{t}(\mathcal{Q},\mathcal{D}_{2}),\dots,f_{t}(\mathcal{Q},\mathcal{D}_{k})]$
$\mathbf{z}_{s}$ $=[f_{s}(\mathcal{Q},\mathcal{D}_{1}),f_{s}(\mathcal{Q},\mathcal{D}_{2}),\dots,f_{s}(\mathcal{Q},\mathcal{D}_{k})]$
The student model $f_{s}$ is trained to mimic the teacher’s output distribution over the candidate texts by minimizing the Kullback-Leibler (KL) divergence between the two softened probability distributions:
$\mathcal{L}_{\text{KD}}=D_{\mathrm{KL}}\left(\mathrm{softmax}\left(\frac{\mathbf{z}_{t}}{T}\right)\bigg|\mathrm{softmax}\left(\frac{\mathbf{z}_{s}}{T}\right)\right)$
where $T$ is the temperature hyperparameter. A higher temperature creates a softer probability distribution, which can help in transferring more nuanced information from the teacher. KD thus provides an essential bridge, allowing the efficiency of bi-encoders to approach the effectiveness ceiling set by cross-encoders.
6.2 Deep Interaction Models: The Cross-Encoder for Reranking